Please be clear with the steps of the solution

roblem (4: a- A Tayyar company sells Motor casings to other manufacturers, would like to uce its inventory cost by determining the optimal number of Motor casings to obtain per order. The annual demand is 10,000 units, the ordering cost is $ 100 per order and the average carrying cost per unit per year is $ 5.0. 1- Calculate the optimal number of units per order. 2- The relevant total annual inventory cost is the sum of the ordering cost and carrying costs. 3- Discuss the Sensitivity Analysis with the EOQ Model.
media%2F36a%2F36a3e1cc-e953-4f53-b7d8-24
Answer #1

1. annual demand- 10,000 units, ordering cost is $100 per order, and inventory carrying cost per unit is $5

EOQ= 2AO /C , where A is annual demand, O is ordering cost and C is carrying csot

EOQ= \sqrt{2*10,000*100/5} = 632.45 units, take it as 632 units per order

2. number of orders= 10,000/632= 15.822 orders

total ordering cost= 15.822*100= $1582.2

3. sensitivity analysis and EOQ: sensitivity analysis is the study of probabilistic values in a given problem, where as EOQ is the study of fixed unit ordering for a given demand.